### v1

# DEEP LEARNING-DRIVEN MALWARE CLASSIFICATION WITH API CALL SEQUENCE ANALYSIS AND CONCEPT DRIFT HANDLING

Abstract
Malware classification in dynamic environments presents a significant challenge due to concept drift, where the statistical properties of malware data evolve over time, complicating detection efforts. To address this issue, we propose a deep learning framework enhanced with a genetic algorithm to improve malware classification accuracy and adaptability. Our approach incorporates mutation operations and fitness score evaluations within genetic algorithms to continuously refine the deep learning model, ensuring robustness against evolving malware threats. Experimental results demonstrate that this hybrid method significantly enhances classification performance and adaptability, outperforming traditional static models. Our proposed approach offers a promising solution for real-time malware classification in ever-changing cybersecurity landscapes.
Keywords API calls  $\cdot$ Malware Classifier  $\cdot$ $n$-grams  $\cdot$ Portable Executable  $\cdot$ Fitness Score  $\cdot$ Mutation
1 Introduction
In the field of cybersecurity, the examination of Portable Executable (PE) files plays a crucial role in detecting and analyzing malware. This study introduces a novel method for dynamic PE malware analysis, leveraging advanced deep-learning techniques to enhance classification accuracy and efficiency. Our proposed model employs neural networks to categorize malware samples into distinct families based on their behaviour and intrinsic characteristics.
A standout feature of our approach is the integration of Genetic Algorithms to manage concept drift, which is a common challenge in malware detection due to the evolving nature of malicious software. Concept drift refers to the changes in the statistical properties of the target variable, necessitating continuous adaptation of the model to maintain its performance over time. By incorporating Genetic Algorithms, our classification system not only remains effective but also adapts dynamically to new and emerging threats, ensuring robust and reliable malware detection. This methodology underscores the importance of adaptability and resilience in cybersecurity strategies, paving the way for more advanced and responsive defense mechanisms against sophisticated cyber threats.
This paper aims at exploring the application of deep learning techniques with a combination of genetic algorithm approaches for dynamic malware analysis, focusing on $n$-gram API calls. We aim at leveraging deep learning techniques to identify patterns and behaviours in malware samples, enhancing our understanding and classification capabilities with concept drift.
The remaining sections of the chapter are structured as follows: Section 2 explains the basic concepts of concept drift, its types, and the role of genetic algorithms in concept drift handling and sandbox environments. Section 3 discusses the related work on concept drift. Section 4 details the data collection process. Section 5 presents our DL and genetic algorithms-based framework for malware classification, covering preprocessing, NLP, feature selection, malware classification phases, and concept drift handling. Section 6 describes our experimental setup, including obtaining an API key, data extraction, malware classification, and concept drift handling, and provides a detailed analysis of results obtained by using different deep learning algorithms combined with genetic algorithms. Section 7 compares our approach with state-of-the-art techniques in malware detection. Section 8 covers the threads to validity. Finally, Section 9 presents the conclusion and the future work of our paper.
2 Basic Concepts
In this section, we introduce the fundamental concepts and terminologies necessary to understand this paper, focusing on concept drift, genetic algorithms, and sandbox environments.
2.1 Concept-drift
Concept drift [1][2] refers to the phenomenon in which the statistical properties of a target variable, which a model is trying to predict, change over time in unforeseen ways. This can happen for various reasons, such as changes in the environment, the data source, or the relationships between variables. Concept drift can be categorised into several types based on the nature of the change:
  1. 1.
    Sudden Concept Drift: This type of drift occurs when there is an abrupt and significant change in the target variable’s distribution. It can be caused by events such as a sudden change in user behaviour or a change in the underlying process generating the data.
  2. 2.
    Incremental Concept Drift: Incremental drift refers to a gradual, continuous change in the target variable’s distribution over time. This type of drift can be challenging to detect, as the changes are subtle and accumulate over time.
  3. 3.
    Recurring Concept Drift: Recurring drift occurs when the target variable’s distribution follows a recurring pattern or cycle. For example, seasonal changes in user behaviour or periodic changes in the underlying process can lead to recurring drift [3].
  4. 4.
    Concept Drift by Context: This type of drift occurs when the relationship between the input variables and the target variable changes based on the context or conditions under which the data is generated. Contextual drift can be particularly challenging to model, as it requires capturing the contextual information along with the data.
  5. 5.
    Covariate Shift: Covariate shift refers to a change in the distribution of the input variables while the relationship between the input and target variables remains the same. This type of drift can be addressed by reweighting the data to match the new distribution of the input variables [4].
Detecting and adapting to concept drift is essential for maintaining the performance of machine learning models in dynamic environments. Various techniques, such as online learning, ensemble methods, and monitoring performance metrics over time, can be used to mitigate the impact of concept drift.
2.2 Techniques to Handle Concept Drift
Handling concept drift is crucial for maintaining the performance of machine learning models in dynamic environments. Below, we present some several techniques commonly used to address concept drift.
  1. 1.
    Re-training: Periodically retraining the model using the most recent data can help the model adapt to concept drift. This approach is effective for both sudden and incremental drift.
  2. 2.
    Ensemble Methods: Ensemble methods, such as bagging and boosting, can improve model robustness to concept drift by combining the predictions of multiple models trained on different subsets of data or with different algorithms.
  3. 3.
    Online Learning: Online learning techniques allow models to be updated continuously as new data becomes available. This approach is well-suited for handling incremental drift.
  4. 4.
    Change Detection Algorithms: Change detection algorithms, such as Cumulative Sum (CUSUM) and Page-Hinkley, can be used to detect concept drift by monitoring the model’s performance metrics over time.
  5. 5.
    Instance Weighting: Instance weighting techniques, such as Importance Weighted Cross Validation (IWCV), can be used to give more weight to recent data instances, helping the model adapt to incremental drift.
  6. 6.
    Feature Selection and Extraction: Adapting the feature set based on the changing data distribution can help improve the model’s performance in the presence of concept drift.
  7. 7.
    Windowing: Using a sliding window approach to train the model on the most recent data can help the model adapt to gradual changes in data distribution.
  8. 8.
    Dynamic Feature Adaptation: Modifying the feature representation based on the current data distribution can help the model better capture the underlying patterns in the data.
  9. 9.
    Memory-based Methods: Memory-based methods, such as k-nearest neighbours (KNN), can be adapted to store historical data and update the model as new data becomes available, allowing the model to adapt to concept drift.
Each of these techniques has its strengths and weaknesses, and the choice of technique depends on the specific characteristics of the data and the problem at hand.
2.3 Genetic Algorithms for Concept Drift
Genetic algorithms (GAs) [5] can be used to handle concept drift in machine-learning models. Genetic algorithms are optimisation algorithms inspired by the process of natural selection and genetics. They can be applied to adapt the model to changing environments, including concept drift scenarios. Here’s how GAs can be used for concept drift:
  1. 1.
    Feature Selection: Genetic algorithms can be used to dynamically select the most relevant features for the current concept. This can help the model adapt to changing data distributions.
  2. 2.
    Model Selection: GAs can be used to evolve the structure of the model, such as the type of model or the hyperparameters, to better fit the current concept.
  3. 3.
    Ensemble Generation: GAs can be used to create ensembles of models with different structures or parameters, allowing the ensemble to adapt to changing concepts.
  4. 4.
    Hyperparameter Optimization: GAs can be used to optimize the hyperparameters of the model to improve its performance in the current concept.
  5. 5.
    Instance Selection: GAs can be used to select a subset of instances from the historical data that are most relevant to the current concept, helping the model adapt more effectively.
By using genetic algorithms in these ways, machine learning models can be made more robust to concept drift, allowing them to maintain high performance even as the data distribution changes over time.
2.4 Sandbox
A sandbox is a security mechanism used in software development, testing, and cybersecurity to isolate running programs or processes from the rest of the system. Inspired by children’s sandbox play areas, it provides a controlled environment where untrusted or potentially harmful code can be executed without risking the underlying system [6].
2.5 Working Principle of a Sandbox
  1. 1.
    Resource Control: Sandboxes limit the resources available to processes, such as files, network resources, memory, and CPU usage, to mitigate potential damage from malicious code.
  2. 2.
    Monitoring and Analysis: Sandboxes include monitoring capabilities to log and analyze system calls, network activity, file operations, and other interactions to detect suspicious behavior.
  3. 3.
    Dynamic Analysis: Sandboxes use real-time dynamic analysis to detect and respond to threats, identifying attempts to exploit vulnerabilities or execute malicious actions [7].
  4. 4.
    Containment: Upon detecting malicious behavior, sandboxes can terminate the offending process or revert changes to the system to contain the threat.
2.6 Cuckoo Sandbox
Cuckoo Sandbox [^1] https://github.com/cuckoosandbox/ [8] [9] is an open-source automated malware analysis system designed to analyse suspicious files and URLs within a controlled environment. It utilises a virtualised environment to execute potentially malicious code safely and provides detailed analysis reports for further investigation.
2.6.1 System Requirements
Before installing Cuckoo Sandbox, we have to ensure that our system meets the following requirements:
  • Availability of Linux-based operating system (e.g., Ubuntu, Debian, CentOS)
  • Adequate hardware resources (CPU, RAM, Disk space) for running virtual machines
  • Availability of Virtualization software (e.g., VirtualBox, VMware, KVM)
  • Python 3.x installed
3 Related Work
Chen et al. [10] distinguished between feature-space drift and data-space drift in malware detectors, highlighting the predominant influence of data-space drift on model degradation over time. Their findings underscored the necessity for further exploration into the implications of feature-space updates, particularly in the context of Android malware datasets like AndroZoo and EMBER.
Jameel et al. [11] conducted a critical review of concept drift’s adverse effects on machine learning classification models. They proposed the ACNNELM model as optimal for Big Data stream classification but noted the absence of critical parameters for advanced ML models like deep learning. The review also highlighted the lack of a matrix model to measure adaptability factors, suggesting avenues for future research in model evaluation and optimization.
Hashmani et al. [12] presented a systematic literature review on concept drift evolution in machine learning approaches. Their comprehensive synthesis and categorization of existing research provided valuable insights into the state of the art in handling concept drift. However, the paper itself did not contribute new methodologies, serving primarily as a reference for researchers seeking to understand current trends and challenges in the field.
Lu et al. [2] conducted an extensive review focusing on concept drift in machine learning, covering detection, understanding, and adaptation strategies across numerous studies. While offering valuable insights into the breadth of research in this area, the paper did not introduce novel methodologies, functioning primarily as a compilation and analysis of existing approaches.
Farid et al. [13] proposed an adaptive ensemble classifier for mining concept drifting data streams. Their methodology addressed the challenge of concept drift by leveraging an ensemble approach, demonstrating promise for real-world applications with evolving data streams. However, the effectiveness of their approach was contingent upon the selection of appropriate base classifiers, highlighting a potential area for improvement in future research.
4 Data Collection and Preprocessing
In this section, we discuss the process of data collection from VirusShare [14] and VirusTotal [15].
Figure 1: Data Collection Process
  1. 1.
    Downloading Malware Hashes: The process starts with obtaining malware hashes from VirusShare [14], a platform known for its collection of malware samples.
  2. 2.
    Passing Hashes to VirusTotal: These hashes are then sent to VirusTotal [15] using an API key. VirusTotal analyzes files and URLs, stores the hashes, and signatures in their repository to detect malicious content.
  3. 3.
    Downloading Antivirus Scan Results: Subsequently, the results from 70 different antivirus scans are downloaded for each hash from VirusTotal in JSON format.
  4. 4.
    Categorizing Malware: The JSON results are analyzed and categorized into different malware types such as adware, backdoor, trojan, spyware, virus, downloader, and worms. These categories are stored in seven different files initially. The antivirus scan results are stored in JSON format, and after processing, the malware types are documented in XLS format.
  5. 5.
    Passing Results Back to VirusShare’s API: The categorized results are then passed back to VirusShare’s API.
  6. 6.
    Downloading Malware Samples: Based on the categorized hashes, specific malware samples are downloaded from VirusShare.
5 Proposed Architecture
In this section, we discuss the architecture and methodology of our proposed model. The proposed model is shown in Figure 2.
Figure 2: Proposed Architecture for Malware Analysis with Concept Drift handling
The detailed design of the proposed neural network architectures is presented here. Our approach consists of five phases. Below, we explain these phases in detail.
Phase 1: Preprocessing Phase
In the preprocessing phase, we perform the following activities.
  1. 1.
    API Dataset & Cuckoo Sandbox Analysis: The process starts with a dataset that undergoes this analysis.
  2. 2.
    Data Division: The data is divided into training and testing sets, both processed through JSON format.
  3. 3.
    API Elements Extraction: Various API elements like APICategory, APIName, APIArgument, and APIreturn are extracted from the API call sequence as shown in Figure 4.
Phase 2: NLP Phase
In NLP phase, we perform the following activities.
  1. 1.
    Creating $n$-grams and Unique $n$-grams: We extracted $n$-grams from the Cuckoo report (JSON). From these $n$-grams, we created unique unigram, bigram, and trigram corpora for each class of malware and benign samples.
  2. 2.
    $n$-grams after Processing: The following are the examples of unigrams, bigrams, and trigrams that we have used in this paper.
    • Unigram: LdrLoadDll_urlmon_urlmon.dll
    • Bigram: NtAllocateVirtualMemory_na,
      LdrLoadDll_ole32_ole32.dll
    • Trigram: LdrUnloadDll_SHELL32,
      LdrLoadDll_SETUPAPI_SETUPAPI.dll,
      LdrGetProcedureAddress_ole32_OleUninitialize
  3. 3.
    Calculating TF: Term Frequency is applied to transform the text data. It tokenizes text, counts the occurrences of each token, and computes TF weights. These weights reflect the importance of each token in a document relative to the entire corpus.
  4. 4.
    Refining Feature Set: A refined feature set is obtained after filtering based on frequency.
Phase 3: Feature Selection Phase
In the feature selection phase, we perform the following activities.
Figure 3: Mutated Features Creation using Genetic Algorithms
  1. 1.
    Creating API Call Frequency Feature Set: Explore features derived from API call frequency to understand the system behaviour and usage patterns.
    Figure 4: Potential Feature Selection after Cuckoo Analysis
  2. 2.
    Applying Filter Based Feature Selection: Apply filter-based techniques (e.g., mutual information, correlation analysis) to select the most informative features.
  3. 3.
    Refining Feature Set: Eliminate redundant or irrelevant features using set hybrid feature selection techniques to ensure that the final set is discriminative and predictive.
Phase 4: Malware Classification Phase
In the malware classification phase, we perform the following activities.
  1. 1.
    Applying Deep Learning Technique: In this context, deep learning techniques such as ANN, RNN and CNN are used for malware classification. Deep learning techniques involve the use of neural networks with multiple layers to learn complex patterns and representations from data. This approach is well-suited for tasks such as malware classification, where the data may have intricate patterns that are difficult to capture using traditional machine-learning algorithms.
Phase 5: Concept Drift Handling Phase
In the concept drift handling phase, we perform the following activities.
  1. 1.
    Applying Genetic Algorithms on Malware Original Pattern: Genetic algorithms are used to evolve and optimise solutions to a problem, mimicking the process of natural selection. In the context of malware analysis, genetic algorithms can be applied to evolve or mutate features of malware samples to explore different characteristics and improve classification or analysis results. The mutation occurs in the sub-parts of API sequence features of secondary and tertiary, but the primary remains untouched, as shown in Figure 3 and 5. We have generated 101248 new mutants by applying crossover and mutation on the malware’s original pattern (Unigrams).
    Figure 5: Mutated Features
    Table 1: Selecting new features using Fitness Score
    Malware F1 F2 Fn Top 1500 Features
    Adware 541 1067 1354 F1, F5, F9
    Backdoor 1265 1397 1289 F85, F1101, F2077
    Downloader 1753 824 1250 F651, F1742, F222
    Spyware 891 1925 1065 F2089, F437, F901
    Trojan 1744 1129 1163 F1600, F2200, F882
    Worm 1579 1367 1625 F785, F2023, F1488
    Virus 903 1064 1293 F33, F1999, F777
  2. 2.
    Calculation of Fitness Score: The fitness score [16] is calculated as the edit distance from the target string. The formula for calculating the fitness score for an individual in the population is:
    $\text{{Fitness}}(individual)=\sum_{i=1}^{n}1_{a_{i}\neq b_{i}}$ (1)
    Where:
    • $\text{{Fitness}}(individual)$ is the fitness score of the individual.
    • $n$ is the length of the target string.
    • $a_{i}$ is the $i$-th character of the target string.
    • $b_{i}$ is the $i$-th character of the individual.
    This formula calculates the number of positions in the individual where the character does not match the corresponding character in the target string, summing them up to get the total edit distance. We conducted fitness score calculations on a dataset comprising 101,248 new mutants. From this pool, we selected 10,500 mutants, with the top 1500 from each malware category based on their fitness scores as shown in Table 1. These mutants serve as features, adding approximately $\approx$ 1% to our existing feature corpus.
  3. 3.
    Final Feature Selection using Genetic Algorithm: This phase involves the selection of the final feature set, which consists of the original malware set and the mutated features of malware, as well as benign features. This selection process is crucial for ensuring that the features used for analysis or classification are relevant and effective in distinguishing between malware and benign samples.
    Table 2: CNN Model Architecture
    Layer Details
    Input Shape (88972, 1)
    Conv. Layer 1 Filters = 64, Kernel Size = 3, Activation = ReLU
    MaxPooling Layer 1 Pool Size = 2
    Conv. Layer 2 Filters = 32, Kernel Size = 3, Activation = ReLU
    MaxPooling Layer 2 Pool Size = 2
    Flatten Layer Flatten the input to a 1D array
    Dense Layer 1 Neurons = 128, Activation = ReLU
    Dropout Layer 1 Dropout Rate = 0.3
    Dense Layer 2 Neurons = 64, Activation = ReLU
    Dropout Layer 2 Dropout Rate = 0.3
    Output Layer Neurons = 8, Activation = Softmax
    Table 3: RNN Model Architecture
    Layer Details
    Input Shape (88972, 1)
    SimpleRNN Layer Units = 128, Activation = ReLU
    Dropout Layer 1 Dropout Rate = 0.5
    Dense Layer 1 Neurons = 64, Activation = ReLU
    Dropout Layer 2 Dropout Rate = 0.5
    Output Layer Neurons = 8, Activation = Softmax
    Table 4: ANN Model Architecture
    Layer Details
    Input Shape (88972, 1)
    Dense Layer 1 Neurons = 512, Activation = tanh
    Dropout Layer 1 Dropout Rate = 0.4
    Dense Layer 2 Neurons = 256, Activation = tanh
    Dropout Layer 2 Dropout Rate = 0.4
    Dense Layer 3 Neurons = 128, Activation = tanh
    Dropout Layer 3 Dropout Rate = 0.4
    Dense Layer 4 Neurons = 64, Activation = tanh
    Dropout Layer 4 Dropout Rate = 0.4
    Output Layer Neurons = 8, Activation = Softmax
  4. 4.
    Applying Deep Learning Techniques: In this context, the ANN, RNN and CNN architecture used for deep learning are shown in Table 2, 3 and 4. We have used the same ANN, RNN and CNN architecture that is used in Phase 4.
6 Implementation and Results
In this section, we discuss the experiment setup and analysis of the contained result after the experiment.
6.1 Experimental Setup
Our experimental setup aimed at evaluating the effectiveness of machine-learning techniques in malware classification. It consists of the following components:
  1. 1.
    Analysis Environment
    • Host OS: We used Ubuntu 18.04 LTS on a machine with an Intel i7 processor, 8GB RAM, and a 10TB HDD.
  2. 2.
    Cuckoo Sandbox
    • Version: Cuckoo Sandbox 2.0.7 was employed for malware analysis on the Ubuntu host.
  3. 3.
    Malware Samples
    • We collected two lakhs of a diverse set of malware samples representing seven categories: adware, backdoor, downloader, spyware, trojan, virus, and worm.
  4. 4.
    Windows 10 Environment
    • A separate Windows 10 environment was used with an Intel i7 processor, 128GB RAM, and 5TB storage to collect and analyse dynamic analysis reports from Cuckoo Sandbox.
6.2 Experimental Results
Table 5: Dataset 1 (collected during January to June 2023)
No. Types Test Sample Train Sample Total Sample
1 Adware 406 1580 1986
2 Backdoor 123 551 674
3 Downloader 495 2002 2497
4 Spyware 190 756 946
5 Trojan 695 2873 3568
6 Worm 277 1080 1357
7 Virus 500 1892 2392
8 Benign 1724 6910 8634
Total 4410 17644 22054
For the concept drift handling using genetic algorithms, we used a total of 22054 samples, including 4410 test samples and 17644 train samples, across various malware types as well as Benign samples as shown in Table 5 as dataset (Dataset 1). This dataset was collected from VirusShare from January to June 2023 release.
Table 6 and Figure 6 illustrate the performance of the Artificial Neural Network (ANN) model across different epochs during training. Initially, the model exhibits a high loss of 1.1005 and a low accuracy of 0.5595 in the first epoch. However, as the number of epochs increases, the loss steadily decreases, reaching 0.0761 by the 100th epoch, accompanied by a significant increase in accuracy to 0.9799. The validation metrics mirror this trend, with the validation accuracy reaching 0.9735 at the 100th epoch.
In contrast, the Recurrent Neural Network (RNN) model demonstrates different behaviour. As shown in Table 7 and Figure 7, the RNN model maintains a relatively constant and high loss, ranging from 1.9359 in the first epoch to 1.6867 by the 100th epoch, along with a consistent accuracy of 0.3916. Similarly, the validation metrics remain stagnant, with the validation loss at 1.8181 and validation accuracy at 0.3908 throughout training.
The Convolutional Neural Network (CNN) model shows the most promising results among the three models. As depicted in Table 8 and Figure 8, the CNN model starts with a high loss of 7.8877 in the first epoch but experiences a rapid decrease with increasing epochs, reaching 0.0161 by the 100th epoch. The accuracy also improves significantly, starting at 0.7281 and reaching 0.995 by the 100th epoch. The validation metrics follow a similar pattern, with the validation accuracy peaking at 0.9853 by the 100th epoch.
To summarize, we can say that the CNN model outperforms the ANN and RNN models in terms of achieving lower loss and higher accuracy. The ANN model also shows promising results, with a steady improvement in performance with more epochs. However, the RNN model struggles, showing minimal improvement in loss and accuracy throughout training. These results suggest that CNN model is most suitable for this task, followed by ANN model, while RNN model may not be well-suited for this problem, i.e., classifying the malwares with concept drift.
Table 6: Loss and Accuracy for Different Epochs of ANN
No. Epoch Loss Val Loss Accuracy Val Accuracy
1 1 1.1005 0.5596 0.6485 0.8203
2 20 0.1959 0.1678 0.9461 0.9658
3 40 0.1524 0.149 0.9577 0.9694
4 60 0.1096 0.14 0.973 0.9703
5 80 0.0869 0.1258 0.978 0.9723
6 100 0.0761 0.1226 0.9799 0.9735
Figure 6: Epoch vs Loss and Epoch vs Accuracy for ANN
Table 7: Loss and Accuracy for Different Epochs of RNN
No. Epoch Loss Val Loss Accuracy Val Accuracy
1 1 1.9359 1.812 0.3754 0.3908
2 20 1.785 1.7761 0.3916 0.3908
3 40 1.8376 1.8367 0.3916 0.3908
4 60 1.7867 1.7881 0.3916 0.3908
3 80 1.7376 1.7967 0.3916 0.3908
4 100 1.6867 1.8181 0.3916 0.3908
Table 9 shows the performance metrics obtained after applying Artificial Neural Network (ANN), Convolutional Neural Network (CNN), and Recurrent Neural Network (RNN) on a new dataset of malware, downloaded from VirusShare and released between January and June 2023. It displays the maximum accuracy and loss achieved by each model.
Figure 7: Epoch vs Loss and Epoch vs Accuracy for RNN
Table 8: Loss and Accuracy for Different Epochs of CNN
No. Epoch Loss Val Loss Accuracy Val Accuracy
1 1 7.8877 4.6905 0.7281 0.8645
2 20 0.1254 0.1722 0.9691 0.9764
3 40 0.0505 0.1292 0.9875 0.985
4 60 0.0248 0.157 0.9927 0.9866
5 80 0.0312 0.1081 0.9937 0.9859
6 100 0.0161 0.1846 0.995 0.9853
Figure 8: Epoch vs Loss and Epoch vs Accuracy for CNN
Table 9: Performance metrics obtained using ANN, CNN, and RNN on Dataset 1 (Without concept drift handling)
DL Epoch Loss Val Loss Accuracy Validation
Tech. Accuracy
ANN 100 0.0761 0.1226 0.9799 0.9735
CNN 100 0.0161 0.1846 0.9950 0.9853
RNN 100 1.6867 1.8181 0.3916 0.3908
Table 10 shows the dataset (Dataset 2) used for the concept drift handling approach. It contains information about different types of samples in the dataset, including Adware, Backdoor, Downloader, Spyware, Trojans, Worms, Virus, and Benign files. For each type, the table lists the number of test samples, train samples, and the total number of samples. The total number of samples in the dataset is 22,000, with 4,400 test samples and 17,600 train samples. This dataset was collected from VirusShare from the January to May 2020 release.
Table 10: Dataset 2 (collected during January to May 2020)
No. Types Test Sample Train Sample Total Sample
1 Adware 400 1600 2000
2 Backdoor 132 528 660
3 Downloader 480 1920 2400
4 Spyware 180 720 900
5 Trojan 700 2800 3500
6 Worm 280 1120 1400
7 Virus 520 2080 2600
8 Benign 1720 6880 8600
Total 4400 17600 22000
Table 11: Performance metrics obtained using ANN, CNN, and RNN on Dataset 2 (Without concept drift handling)
DL Epoch Loss Val Loss Accuracy Validation
Tech. Accuracy
ANN 100 0.6250 8.0510 0.9112 0.8946
CNN 100 0.1931 0.2864 0.9250 0.9430
RNN 100 2.1200 2.5670 0.3429 0.3552
Table 12: Performance metrics obtained using ANN, CNN, and RNN on Dataset 2 (with concept drift handling)
DL Epoch Loss Val Loss Accuracy Validation
Tech. Accuracy
ANN 100 0.6314 6.8510 0.9343 0.9184
CNN 100 0.1429 0.2422 0.9314 0.9459
RNN 100 1.9150 2.1563 0.3501 0.3558
In contrast, Table 11 presents the performance of these models on an older dataset from VirusShare, released between January and May 2020, highlighting the increase in validation accuracy as well as the decrease in validation loss. Moreover, Table 12 demonstrates the impact of concept drift on the same old dataset (Dataset 2), showing how the models perform when trained on data that has evolved since its original release. These findings suggest that utilising concept drift in the malware dataset can lead to improved testing accuracy and reduced loss, indicating the effectiveness of the concept drift malware classifier.
7 Comparison with Related work
Our work focuses on improving malware classification using NLP-based $n$-gram API sequence coupled with deep learning and concept drift handling with genetic algorithms. Since we utilise a unique dataset, we lack direct comparisons with existing state-of-the-art techniques. Our approach harnesses the power of genetic algorithms, deep learning and $n$-gram analysis, offering a distinctive perspective on malware detection that can handle concept drift. In the absence of any directly related work, we compare our work with some closely related work.
García et al. [17] proposed a technique to assess the effectiveness of transfer learning (TL) methods for malware classification in the presence of concept drift, focusing on various time periods and learning scenarios. The study utilised five TL algorithms—TrAda, CORAL, DAE, DTS, and TIT— and evaluated their effectiveness in handling concept drift in malware classification. TrAda, CORAL, and DAE were identified as the most effective TL algorithms, consistently achieving Matthews correlation coefficients (MCC) greater than 0.9. Among the machine learning (ML) algorithms, Random Forest (RF) demonstrated competitive performance, especially in the inductive TL setting. The dataset used in the study was not explicitly disclosed by the authors, making the exact source of the malware and benign dataset unclear.
Table 13: Quantitative comparison of some deep learning-based malware detection and classification techniques with concept drift
S No. Authors $n$-gram API Concept Drift Dataset Source Dataset Features
Seq. used Tech. used Size Used
1 García et al. [17] Transfer learning Not Disclosed 9196 APIs, signatures,
and network
2 Fernando et al. [18] Genetic Algorithms Elderan dataset 2720 API Calls
3 Proposed Work Genetic Algorithms VirusShare 22054 API Calls
Fernando and Komninos [18] introduced the feSAD ransomware detection framework, which leverages machine learning to adapt to concept drift. The framework aims at enhancing ransomware detection rates by calibrating drift thresholds and identifying abnormal drift samples. Compatible with various machine learning algorithms, it has demonstrated effectiveness in detecting ransomware amidst concept drift and can be tailored for different malware types. Experimental results, including 720 ransomware samples and 2000 benign samples from the Elderan dataset, show high detection rates and precision, especially with the Random Forest algorithm. Komninos highlighted the Random Forest’s stability in ransomware detection. However, they noted that in test batch 2, a high volume of ransomware samples caused abnormal concept drift, leading to reduced detection rates and statistical drift, suggesting system retraining to avoid detection decline.
8 Threats to Validity
There are some potential threats to the validity of the proposed model and its results. We discussed them below.
Internal Validity
  • Algorithm Performance: The effectiveness and the performance of the genetic algorithm approach for concept drift handling could be influenced by the specific parameters and configurations chosen, which may impact the results.
External Validity
  • Concept Drift Representation: The concept drift handling approach may not fully capture the complexity and dynamics of concept drift in real-world malware datasets, affecting the effectiveness of the proposed approach in practical scenarios.
Construct Validity
  • Model Architecture: The specific architecture of the neural network used for classification may not be optimal for handling concept drift or may not fully leverage the additional features added through genetic algorithms.
Conclusion Validity
  • Evaluation Metrics: The evaluation metrics used (e.g., accuracy, loss) may not fully capture the effectiveness of the approach in handling concept drift and distinguishing between malware and benign samples, potentially leading to biased conclusions.
Addressing these threats involves conducting thorough experiments with diverse datasets, carefully selecting parameters and configurations for the genetic algorithm and neural network, and considering the implications of concept drift on the model’s performance and generalizability.
9 Conclusion and Future Work
In this paper, we presented a deep learning approach for dynamic PE malware analysis, focusing on handling concept drift. Our model includes the phases of preprocessing, NLP processing, feature selection, malware classification, and concept drift handling.
We used API call frequency features for initial selection and added 10,500 features using genetic algorithms, improving the model’s ability to distinguish malware from benign samples and handle concept drift.
Our neural network has an input layer with 88,972 neurons, three hidden layers, and an output layer with eight neurons for multi-class classification. ReLU activation and Adam optimizer were used for training, resulting in high accuracy and low loss.
We calculated fitness scores for 101,248 new mutants, selecting the top 10,500 for feature augmentation, increasing the feature corpus by approximately 1%.
Experimental results showed significant improvements in classification accuracy and loss reduction across a number of epochs and datasets, indicating the potential of our approach for enhancing malware detection and classification in dynamic PE environments, with concept drift handling.
Scope for Future Research
Below we present some future scope of our research work.
  • Explore the application of other deep learning approaches, such as Long Short Term Memory Networks (LSTMs) and Generative Adversarial Networks (GANs), to broaden the scope of the research and potentially uncover new insights.
  • Investigate the use of genetic algorithm approaches for feature selection and feature creation to enhance the robustness and accuracy of the models.
  • Experiment with transfer learning techniques to leverage pre-trained models for malware classification tasks, potentially improving performance and reducing training time.
  • Conduct cross-domain analysis by applying the developed models to different types of malware datasets, assessing their generalizability and adaptability to diverse malware threats.
10 Code and Dataset Availability
The research code is available on GitHub at the following link: https://github.com/bishwajitprasadgond/MalClassCD. For access to the dataset used in this research, please send a request via email to bishwajitprasadgond@gmail.com.
11 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.